Representation of intergers

Change the following number systems to base-10 numbers.

  • 11011010012
  • 55228
  • 12CF16
1
Expert's answer
2022-05-19T17:52:14-0400

a.


"(1101101001)_2=2^9+2^8+2^6+2^5+2^3+2^0"

"=512+256+64+32+8+1=873"

b.


"(5522)_8=5(8)^3+5(8)^2+2(8)^1+2(8)^0"

"=2560+320+16+2=2898"

c.


"(12CF)_{16}=1(16)^3+2(16)^2+12(16)^1+15(16)^0"

"=4096+512+192+15=4815"